When a 100-m metal pipe is heated from 20°C to 43.5°C, it lengthens by 4.00 cm. What type of pipe is this?
brass
aluminum
gold
copper
1
Expert's answer
2013-01-14T09:23:23-0500
Relative expansion dL/L isequal to alpha*dT, where alpha is the coefficient of linear expansion and dT is the temperature increment.
We have
L = 100 m
dl = 4 cm,
dT = (43.5 - 20) °C = 23.5 °C

Thus, we can find
alpha = dL/(L*dT) = 4*10^(-2) m / (100 m * 23.5 °C) = 1.702 * 10^(-5) 1/°C.

By referring to a table of coefficients for different metals (e.g. https://en.wikipedia.org/wiki/Thermal_expansion) we find that it's closest to that of copper.

Answer: copper